OpenStudy (twizttiez):

What is the value of the function y=2x+3 when x = –1? A. 5 B. –5 C. 2 D. 1

OpenStudy (twizttiez):

@Reaper534

OpenStudy (anonymous):

easy plug in -1 for x y=2(-1)+3 y=-2+3 y=?

OpenStudy (twizttiez):

1

OpenStudy (twizttiez):

So D right?

OpenStudy (anonymous):

good

OpenStudy (twizttiez):

Thanks i have 4 more could you help me through them?

OpenStudy (anonymous):

sure

OpenStudy (twizttiez):

Ok thanks one sec